OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 4 and a6 = 39 ? 209 213 217 221

OpenStudy (anonymous):

The common difference is d = (a6 - a1)/5 = (39 - 4)/5 = 7 Then a32 = a1 + 31*7 = 4 + 217 = 221

OpenStudy (anonymous):

danke!